OpenStudy (jazzyfa30):

costheta=4/5 and 270Degrees

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

direct formula this time - no reference triangle business

ganeshie8 (ganeshie8):

use below : \(\cos 2 \theta = 2\cos^2 \theta - 1\)

OpenStudy (jazzyfa30):

cos2theta =2(4/5)^2-1

ganeshie8 (ganeshie8):

yup ! simplify

OpenStudy (jazzyfa30):

7/25

ganeshie8 (ganeshie8):

\(\color{red}{\checkmark}\)

OpenStudy (jazzyfa30):

i need help with 5 more trig problems is that ok

ganeshie8 (ganeshie8):

yea sure